Herschel Status Report - February 2010
24 Feb 2010 10:53
Report for period 22 January to 19 February 2010
Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected.
The instrument and Instrument Control Centre (ICC) reviews for both PACS and SPIRE have been held. These reviews have determined the readiness of the two instruments for Routine Science Phase operations and for the upcoming first open call for observing proposals. The still outstanding HIFI and "Herschel observatory level" reviews, to declare readiness for issuing the first Open Time AO (planned on 20 May 2010) by ESA's Director of Science and Robotic Exploration, will be held at the beginning of March 2010.
The spacecraft is operating nominally.
Star tracker warm pixels
Further results from more than 100 dedicated PACS pointings show no evidence of any serious deviation in the Absolute Pointing Error from a Gaussian distribution with a 1-sigma error of just under 2 arcseconds. This is still within specifications and in line with pre-flight predictions.
It is now understood why these "warm" pixels feature so prominently in distorting large spatial maps of up to 2 × 2 square degrees: unlike the case of stable pointings a large fraction of the star tracker CCD is used in producing maps and thus it is likely to encounter blemishes more frequently. Scans that cause guide stars to cross blemished star tracker CCD pixels for tens of seconds in scanning maps can easily lead to overall (time-integrated) distortions that far exceed what is expected for fixed pointings which are affected by at most a few hundred pixels out of a quarter million. Under the circumstances, reduction of the CCD temperature from +13 °C to -10 °C will be accommodated in the normal mission planning process and it is expected to complete the exercise by end-March 2010. This allows increasingly utilising "single spatial pixel on the sky" spectroscopy observations as HIFI is brought back into operations.
Despite this reassuring news the Herschel operation teams remain vigilant concerning the appearance of new "warm" pixels. These have recently been reported by industry to appear at a frequency of one per day at the current operating temperature of 13 °C. Full CCD dumps will be produced every 2 weeks for monitoring purposes. MOC/ Flight Dynamics have made an accepted proposal as to how these pixels can be masked, though it is hoped that final cooling will render them unimportant operationally.
Both PACS and SPIRE continue to operate nominally. The two instruments are conducting Routine Science Phase observations with the remaining Science Demonstration Phase observations interleaved. For SPIRE the validation of a few sub-modes for astronomical observation templates (AOTs) is also performed. This affects less than 10% of SPIRE observations.
The full recovery of HIFI is well on track.
Preparations are progressing for the opening of the online Herschel Science Archive to the scientific community. This archive will contain publicly available observations from the Science Demonstration Phase, as well as selected observations from the Performance Verification Phase. The Herschel Interactive Data Processing Environment (HIPE) software will then also be distributed to a larger community for the first time.